OpenStudy (anonymous):

What is the solution of this system of nonlinear equations? (1) x2 + y2 = 25 (2) y = x2 + 6

OpenStudy (anonymous):

(0, 5) (1, 5) and (–1, 5) (5, 0) and (–5, 0) no solution

OpenStudy (anonymous):

@amistre64

OpenStudy (owlcoffee):

since equation 2 already has a variable solved for: \[y=x ^2 +6\] All we have to do is plug it on the first one: \[x^2+(x^2+6)^2=25\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so its no solution

OpenStudy (anonymous):

x^4+13 x^2+36 = 25

OpenStudy (amistre64):

id simply test the options to see which one makes both equation the same

OpenStudy (amistre64):

(1) x2 + y2 = 25 (2) y = x2 + 6 when x=0, y does not equal 5 for both of them when y=0, x does not equal 5 for both of them test out the remaining point ... if it fails then we know its D or not